OpenStudy (anonymous):

what is the solution to the compound inequality? -8<2x-4<2

OpenStudy (anonymous):

-8<2x-4 -4<2x -2<x 2x-4<2 2x<6 x<3 so answer is -2<x<3

OpenStudy (anonymous):

THANK YOU SOOOOO MUCH!

OpenStudy (anonymous):

Can you help me with another one?

OpenStudy (anonymous):

what is the other one?

OpenStudy (anonymous):

What is the solution to the compound inequality |3x-4| = 5

OpenStudy (anonymous):

the answer of 3x-4 can be -5 or 5 because of absolute valve. 3x-4=5 3x=9 x=3 3 is the first answer. 3x-4=-5 3x=-1 x=-1/3 second answer is -1/3.

OpenStudy (anonymous):

THANK YOU!!!!!!!! ok ok ok can you help me with one more. THIS IS THE LAST ONE!!!

OpenStudy (anonymous):

Which represents the solution to the absolute value equation? 3|2x 4|-1=11

OpenStudy (anonymous):

Did you mean 2x-4 or 2x+4?

OpenStudy (anonymous):

2x+4

OpenStudy (anonymous):

OK. \[3\left| 2x+4 \right|=12\]\[\left| 2x+4 \right|=4\] the answer of 2x+4 can be -4 or 4 because of absolute valve. you should solve these equations to find 2 answers. 2x+4=4 2x+4=-4

OpenStudy (anonymous):

Thank you soooooooooooooooooooooooooooo much!!111!!!